OpenStudy (4meisu):

1/9^3x+1 = 27, solve for x

OpenStudy (anonymous):

no problem I solve it please wait

OpenStudy (anonymous):

\[(19/10)^{3x} \times (19/10)^{1}=27\]

OpenStudy (4meisu):

Would it be \[1/9^{3x+1} = 3^3\] \[3x + 1 = 3\] 3x - 2 x = 2/3 ?

OpenStudy (4meisu):

When you did -6x-1 = 3, you put 2 - 6x Wouldn't it be 4 = 6x? There's a negative 1 so you'd have to add it to the 3.

OpenStudy (anonymous):